OpenStudy (anonymous):

The midpoint of UV is (5, -11). The coordinates of one endpoint are U(3,5). Find the coordinates of endpoint V. Help please! I need this explained, my teacher keeps asking questions like this but I never really learned it. if you explain this one to me I can probably figure out the other ones! thanks <3

jimthompson5910 (jim_thompson5910):

Let V be the point (x,y)

jimthompson5910 (jim_thompson5910):

one endpoint is (3,5), the other is (x,y) to find the x coordinate of the midpoint, you add up the x coordinates and divide by 2 so you do this (x1 + x2)/2 (3+x)/2 then you set it equal to 5 since this is the x coordinate of the given midpoint (5,-11) like this (3+x)/2 = 5 now solve for x (3+x)/2 = 5 3+x = 5*2 3+x = 10 x = 10-3 x = 7

jimthompson5910 (jim_thompson5910):

the y coordinate of point V is found in a similar way, just using the y coordinates

OpenStudy (anonymous):

Alright, thank you a lot c: it really cleared it up for me. I'll try to figure out y for myself.

OpenStudy (anonymous):

so I did for y, but I'm not sure if I did it right. (5+y)/2=-11 (5+y)=-22 y= -27 is this right? I'm always confused when negatives are involved. @jim_thompson5910